OpenStudy (anonymous):

WILL MEDAL For which operations is the set {–1, 0, 1} closed? Choose all answers that are correct. A. addition B. division C. multiplication D. subtraction

OpenStudy (anonymous):

A) what is \(1+1\)?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

but there is more then one ansewr

OpenStudy (anonymous):

evidently it is not closed under addition what about \(\frac{1}{0}\)

OpenStudy (anonymous):

i don't get it can you make it alittle more clear?

OpenStudy (anonymous):

closed means if you perform the operation you get something that is still in the set it is not closed under addition because 1 + 1 = 2 and 2 is not in your set

OpenStudy (anonymous):

it is not closed under division because \(\frac{1}{0}\) is not even a number let alone one of the three numbers in the set

OpenStudy (anonymous):

??? im still confused

OpenStudy (anonymous):

it is not closed under subtraction because \(-1-1=-2\) again not in the set

OpenStudy (anonymous):

what is a set?

OpenStudy (anonymous):

@ryguy717

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thats disappointing

OpenStudy (anonymous):

ok i will

OpenStudy (anonymous):

@princed

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

ANYBODY

OpenStudy (anonymous):

i think its division

OpenStudy (anonymous):

i guess

geerky42 (geerky42):

None. 1/0 = undefinied, which is not part of set {-1, 0, 1}

OpenStudy (anonymous):

thx any other guesses there is more then one anserws

OpenStudy (anonymous):

lol i can't spell XD

geerky42 (geerky42):

Addition won't works because 1+1 = 2. 2 is not part of set {-1, 0, 1}. Subtraction; same reason; 1 - -1 = 2 Division. What I said. So that left us multiplication.

OpenStudy (anonymous):

so division and multiplication? @geerky42

geerky42 (geerky42):

Is 1/0 a part of set {-1, 0, 1}?

OpenStudy (anonymous):

idk

geerky42 (geerky42):

Do you see 1/0 in set {-1, 0, 1}?

OpenStudy (anonymous):

XD im not that smart

OpenStudy (anonymous):

ummm

geerky42 (geerky42):

Ok, this set have three elements: -1, 0, and 1, right? So is there 1/0 in this set?

OpenStudy (anonymous):

umm.... yes?

geerky42 (geerky42):

Really, where? All I see are -1, 0, and 1.

OpenStudy (anonymous):

oh then no. as i said, i'm not that smart

geerky42 (geerky42):

Ok, since in division, you can divide 1 by 0, lead us to 1/0. 1/0 is not in {-1, 0, 1}, so this set is not closed under division. So that left us multiplication. Answer is multiplication, that's it.

OpenStudy (anonymous):

so not divison

geerky42 (geerky42):

No.

OpenStudy (anonymous):

no, as in yes there is or no as in yes there isn't?

geerky42 (geerky42):

No division.

geerky42 (geerky42):

Answer is multiplication. That's it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no other?

geerky42 (geerky42):

None, just multiplication.

OpenStudy (anonymous):

ok thx